### v2

# Orthogonal Finetuning Made Scalable

Abstract
[…]
Orthogonal Finetuning Made Scalable
Zeju Qiu1,†    Weiyang Liu1,2,†,*    Adrian Weller3,4    Bernhard Schölkopf1 1Max Planck Institute for Intelligent Systems    2The Chinese University of Hong Kong 3University of Cambridge    4The Alan Turing Institute    Equal contribution *Project lead, Correspondence to wyliu@cse.cuhk.edu.hk spherelab.ai/oftv2
1 Introduction
As foundation models continue to improve in performance, recent years have witnessed a paradigm shift from end-to-end learning to a pretraining-finetuning framework. This shift underscores the need for finetuning methods that are both effective and scalable. Owing to its training stability and adaptation efficiency, orthogonal finetuning (OFT) Qiu et al. (2023); Liu et al. (2024) has emerged as a promising approach for adapting foundation models to downstream tasks. However, while performing well, OFT incurs high computational and memory costs, limiting its scalability. Motivated by these challenges, we seek to make OFT more scalable to large foundation models.
Figure 1: OFTv2 significantly reduces training time and GPU memory usage without sacrificing performance. The finetuning is performed with Qwen2.5-7B.
Towards this goal, we begin by identifying the key bottleneck that limits OFT’s scalability. At its core, OFT learns layer-shared orthogonal matrices to transform pretrained weight matrices, resulting in a naive weight-centric implementation where forward inference is performed after merging the learned orthogonal matrices into weight matrices during training. The weight-centric implementation thus involves matrix-matrix multiplications with cubic complexity. As weight matrices grow large, this cubic scaling severely limits OFT’s applicability to large foundation models. However, these matrix-matrix multiplications are not fundamentally necessary. We draw inspiration from matrix-free methods Chen (2005), such as the power method and the Lanczos algorithm, which avoid explicit matrix-matrix operations by treating matrices as linear operators applied to vectors. These methods operate entirely through matrix-vector multiplications, applying a matrix to vectors in the appropriate space without ever forming full matrix products. Guided by the same insight, we introduce an input-centric implementation of OFT, in which the learned orthogonal transformations are applied directly to the input vectors during each forward pass, rather than being merged into the weight matrix. This reformulation reduces the complexity from cubic to quadratic. We refer to this new formulation as OFTv2. Despite its simplicity, this change significantly enhances the scalability of OFT, making it suitable for finetuning large foundation models that the original OFT could not handle due to memory constraints.
Another scalability bottleneck in OFT arises from the Cayley parameterization used by Liu et al. (2021a); Qiu et al. (2023); Liu et al. (2024) to preserve orthogonality. While effective, this parameterization involves computing a matrix inverse, which becomes increasingly costly and less numerically stable as weight matrices get larger. To address this, we use a numerically stable yet efficient approximation – the Cayley–Neumann parameterization (CNP) Qiu et al. (2025). By replacing the matrix inverse in the original Cayley transform with a truncated Neumann series, CNP offers improved numerical stability and lower computational cost, particularly in settings where OFT is applied to finetune large foundation models. With CNP, OFTv2 becomes even more scalable and readily applicable for efficient adaptation of such models. In Figure 1, we compare OFT and OFTv2 by performing finetuning tasks on Qwen2.5-7B, which is the largest model that the original OFT can finetune within a single Nvidia H100 (80GB). These empirical results demonstrate that OFTv2 achieves substantial GPU memory savings and training speed-up over the original OFT formulation Qiu et al. (2023).
In practice, finetuning ultra-large foundation models (e.g., LLaMA 3.1-70B Grattafiori et al. (2024), Qwen 2.5-72B Yang et al. (2024a)) typically requires quantization to fit within GPU memory limits. To support this, we follow the general design of the QLoRA framework Dettmers et al. (2023) but replace LoRA with OFTv2. Our input-centric implementation of orthogonal finetuning enables a seamless application to the finetuning of quantized foundation models, resulting in QOFT–an efficient orthogonal finetuning that enables efficient adaptation of quantized ultra-large models. Our major contributions are summarized below:
[…]
  • We apply the Cayley–Neumann parameterization Qiu et al. (2025) in OFTv2. It approximates the Cayley transform with a truncated Neumann series and eliminates matrix inversions.
  • Owing to the new input-centric formulation, we adapt OFTv2 to finetuning quantized foundation models. This enables memory-efficient finetuning of ultra-large models.
  • We apply OFTv2 and its quantized variant to different foundation models (including large language models and text-to-image generative models) across various model scales.
2 Related Work
Parameter-efficient finetuning (PEFT). As foundation models become increasingly large and powerful, there has been growing interest in finetuning them for downstream tasks in a parameter-efficient manner Houlsby et al. (2019); Aghajanyan et al. (2020); Hu et al. (2022a); Edalati et al. (2022); Wang et al. (2022); Gheini et al. (2021); Zaken et al. (2022); Guo et al. (2020); Sung et al. (2021); Ansell et al. (2022); Lester et al. (2021); Li and Liang (2021); Vu et al. (2022); He et al. (2021); Mao et al. (2021); Karimi Mahabadi et al. (2021); Liu et al. (2022); Sung et al. (2022); Chen et al. (2023); Jia et al. (2022); Chen et al. (2022); Zhang et al. (2022); Jie and Deng (2023); Lian et al. (2022); Luo et al. (2023); Zhang et al. (2024); Wu et al. (2024). In particular, reparameterization-based methods (e.g., Aghajanyan et al. (2020); Hu et al. (2022a); Edalati et al. (2022); Zi et al. (2023); Chavan et al. (2023)) are enjoying wide adoption. LoRA Hu et al. (2022a) learns a pair of small low-rank matrices whose product is added to each weight matrix, enabling task adaptation with a small number of trainable parameters. Building on LoRA, several works dynamically adjust the rank across layers to better balance the parameter budget Zhang et al. (2023b); Valipour et al. (2022); Zhang et al. (2023a, 2024). To improve scalability, QLoRA Dettmers et al. (2023) quantizes the frozen base model to 4-bit NormalFloat with double quantization and back-propagates only through LoRA, achieving near full-precision accuracy while drastically lowering memory usage.
Orthogonal Finetuning. Qiu et al. (2023); Liu et al. (2024) propose a reparameterization-based method that learns layer-shared orthogonal matrices to transform neurons, yielding strong generalization and stable training. The is motivated by the observation that hyperspherical energy (i.e., a geometric characterization of neurons on the unit sphere) influences generalization Liu et al. (2018, 2021b); Lin et al. (2020); Liu et al. (2023), and that orthogonal transformations keep this energy invariant Liu et al. (2021a). A growing body of research Ma et al. (2024); Yang et al. (2024b); Gorbunov et al. (2024); Yuan et al. (2024); Feng et al. (2025); Raj and Coyle (2025); Lingam et al. (2024); Bini et al. (2024); Su et al. (2024); Liao and Monz (2024) builds upon the core idea of OFT. Figure 2 provides a comparison between OFT and LoRA. OFT achieves parameter efficiency through sparsity, whereas LoRA relies on a low-rank structure.
Figure 2: Comparison between LoRA and OFT.
3 OFTv2: Faster and More Scalable
3.1 Preliminaries
Let $\bm{W}=[\bm{w}_{1},\cdots,\bm{w}_{n}]\in\mathbb{R}^{d\times n}$ be a weight matrix with columns $\bm{w}_{i}\in\mathbb{R}^{d}$. In a linear layer, the forward pass is $\bm{z}=\bm{W}\bm{x}$, where $\bm{x}\in\mathbb{R}^{d}$ is the input and $\bm{z}\in\mathbb{R}^{n}$ is the output. OFT reparameterizes the weight matrix with $\bm{W}_{\text{OFT}}=\bm{R}\bm{W}_{0}$ where $\bm{W}_{0}$ is the pretrained weight matrix and $\bm{R}\in\mathbf{R}^{d\times d}$ is an orthogonal matrix. OFT only learns $\bm{R}$ for adapting the pretrained model to downstream tasks. To enforce orthogonality, Liu et al. (2021b); Qiu et al. (2023); Liu et al. (2024) parameterize $\bm{R}$ using the Cayley transform: $\bm{R}=(\bm{I}+\bm{Q})(\bm{I}-\bm{Q})^{-1}$, where $\bm{Q}$ is a skew-symmetric matrix satisfying $\bm{Q}=-\bm{Q}^{\top}$. To further improve parameter-efficiency, OFT constrains the orthogonal matrix $\bm{R}$ to have a block-diagonal structure: $\bm{R}=\text{Diag}(\bm{R}_{1},\cdots,\bm{R}_{r})$ where for any $i$, $\bm{R}_{i}\in\mathbb{R}^{b\times b}$ is a small orthogonal matrix and $b\cdot r=d$. Each $\bm{R}_{i}$ can be parameterized using the Cayley transform. This block-diagonal form imposes a sparsity pattern on $\bm{R}$, effectively making it a sparse orthogonal matrix. Leveraging this structure, Liu et al. (2024) further enhances parameter efficiency using butterfly factorization.
3.2 From Weight-centric Implementation to Input-centric Implementation
OFT performs finetuning by learning an orthogonal matrix to directly transform the weight matrix, which naturally leads to a weight-centric implementation of the forward pass:
[…]
The original OFT first performs a weight transform by computing $\bm{W}_{\text{OFT}}^{\top}=\bm{W}_{0}^{\top}\bm{R}^{\top}$ (i.e., a matrix-matrix multiplication) and then computes the results of a linear layer with the equivalent weight matrix $\bm{W}_{\text{OFT}}^{\top}$ (i.e., a matrix-vector multiplication). This incurs $\mathcal{O}(nd^{2})$ complexity due to the matrix-matrix multiplication. Inspired by matrix-free methods for solving linear systems, we observe that OFT’s forward pass can be interpreted as two linear maps applied to the input. This leads to an input-centric implementation
[…]
where only two matrix-vector multiplications are required, reducing the complexity from cubic to quadratic: $\mathcal{O}(nd+d^{2})$. This simple conceptual shift in implementation entails a substantial speed-up in training time and reduction in GPU memory.
3.3 Approximate Orthogonality via Cayley-Neumann Parameterization
The Cayley parameterization constructs an orthogonal matrix $\bm{R}$ with $(\bm{I}+\bm{Q})(\bm{I}-\bm{Q})^{-1}$, where $\bm{Q}$ is a skew-symmetric matrix. One limitation of this formulation is that it only generates rotation matrices, though empirical studies Liu et al. (2021a); Qiu et al. (2023); Liu et al. (2024) suggest that this restriction does not negatively affect performance. More critically, computing a matrix inverse introduces numerical instability and additional computational overhead, making it challenging to scale to large orthogonal matrices. To address this, we use the Cayley-Neumann parameterization proposed by Qiu et al. (2025), where the matrix inverse is approximated by a truncated Neumann series:
$\bm{R}$ $=(\bm{I}+\bm{Q})(\bm{I}-\bm{Q})^{-1}=(\bm{I}+\bm{Q})\big(\sum_{i=0}^{\infty}\bm{Q}^{i}\big)$
$\approx(\bm{I}+\bm{Q})\big(\bm{I}+\sum_{i=1}^{k}\bm{Q}^{i}\big),$
where larger $k$ leads to better approximation. Removing the matrix inversion improves training stability. The Neumann series approximation converges in the operator norm if $\|\bm{Q}\|<1$. This condition is naturally satisfied in practice: to start from the pretrained model, OFT initializes the orthogonal matrix $\bm{R}$ as the identity, which requires $\bm{Q}$ to start as a zero matrix. Since finetuning begins with a small learning rate and typically involves relatively few steps, $\bm{Q}$ tends not to drift far from zero. Empirically, even if $\|\bm{Q}\|$ slightly exceeds $1$, it does not harm OFT’s training stability, as we use only a finite number of Neumann terms.
Custom CUDA kernel for skew-symmetric matrices. To maximize GPU memory efficiency, we leverage the skew-symmetric structure of $\bm{Q}\in\mathbb{R}^{n\times n}$, where $Q_{ii}=0$, $Q_{ij}=-Q_{ji}$. By storing only the upper triangular part as a vector, we reduce the storage requirement from $n^{2}$ to $\frac{n(n-1)}{2}$. During the forward pass, $\bm{Q}$ is reconstructed on-the-fly using a highly optimized custom CUDA kernel that significantly accelerates this process.
4 QOFT: Adapting OFTv2 to Finetuning Quantized Foundation Models
While PEFT methods primarily aim to reduce optimizer memory by minimizing trainable parameters, the growing scale of foundation models has shifted the memory bottleneck to the pretrained weights themselves. As model dimensions grow, these frozen parameters increasingly dominate memory consumption during training Kim et al. (2023). To address this emerging challenge, we argue that truly scalable OFT must operate directly on quantized model representations, such as NormalFloat4 Dettmers et al. (2023) and AWQ Lin et al. (2024). This represents a critical shift that enables OFT to scale effectively.
To this end, we introduce QOFT, a natural extension of OFTv2 for quantized foundation models. QOFT largely follows the framework of QLoRA Dettmers et al. (2023). Specifically, the quantized low-bit weight matrices are first dequantized to higher precision, after which the parameter-efficient adaptation is carried out in the higher-precision space. Formally, the forward pass of QOFT can be written as
$\bm{z}=\underbrace{\text{Dequant}(\bm{W}_{\text{quant}})^{\top}}_{\text{Fronzen}}\underbrace{\bm{R}^{\top}}_{\text{Trainable}}\bm{x}$ (3)
The update of OFTv2’s orthogonal matrix $\bm{R}$ is performed in high precision (e.g., BF16). We denote the dequantization function as $\text{Dequant}(\cdot)$ and follow QLoRA’s design by adopting a double quantization strategy, where the quantization parameters of the weight matrices are themselves quantized to further reduce GPU memory usage.
Flexible quantized finetuning via OFTv2. We now explain why the weight-centric implementation of OFT is ill-suited for quantized foundation models. Computing the matrix product $\bm{W}_{\text{quant}}^{\top}\bm{R}^{\top}$ involves rotating (or reflecting) a quantized weight matrix, which requires first dequantizing it to higher precision before applying the transformation. While this is mathematically valid, it makes OFT dependent on the specific quantization method used. Different quantization schemes may require different treatments for computing $\text{Dequant}(\bm{W}_{\text{quant}})^{\top}\bm{R}^{\top}$, introducing unnecessary complexity. In contrast, the input-centric implementation avoids this issue by fully decoupling OFT from weight quantization. It applies the learned orthogonal matrix $\bm{R}^{\top}$ to the input $\bm{x}$. The subsequent forward pass proceeds as usual under any quantization strategy. As a result, OFTv2 becomes a quantization-agnostic PEFT method compatible with arbitrary weight quantization schemes.
QOFT vs. QLoRA. We now look into the forward pass of QLoRA: $\bm{z}=\text{Dequant}(\bm{W}_{\text{quant}})^{\top}\bm{x}+(\bm{A}\bm{B})^{\top}\bm{x}$ where $\bm{A}\in\mathbb{R}^{d\times r}$ and $\bm{B}\in\mathbb{R}^{r\times n}$ are low-rank matrices and $r\ll\min(d,n)$ is usually quite small. First, QOFT is more suitable for post-training quantization when merging the finetuned weights back into the quantized model. In QLoRA, the equivalent weight $\bm{W}+\bm{A}\bm{B}$ can alter the dynamic range (i.e., the possible minimum and maximum values) of the weight matrix, potentially complicating requantization. In contrast, the equivalent weight in QOFT, $\bm{R}\bm{W}$, preserve the dynamic range of individual elements. The worse-case requantization error for QLoRA is always larger than QOFT by $\|\bm{A}\bm{B}\|_{\infty}$. This advantage is also partially supported by recent evidence Tseng et al. (2024); Ashkboos et al. (2024) suggesting that orthogonal transformations can homogenize weight magnitudes and suppress outliers.
Another practical limitation of QLoRA is its training instability. Across various experiments, we observe that QLoRA is prone to loss divergence and unstable optimization. We suspect this arises from the inherently noisier gradients in QLoRA, which adversely affect the finetuned weights. In contrast, QOFT benefits from the orthogonality of $\bm{R}$, which also regularizes the back-propagated gradients. As a result, the adaptation weights in QOFT are better conditioned, and when merged into the pretrained model, they yield a more stable finetuned model. This observation is supported by prior work Qiu et al. (2023); Liu et al. (2024) showing that OFT significantly improves training stability and mitigates catastrophic forgetting.
Figure 3: Comparison between sequential (e.g., OFT) and parallel (e.g., LoRA) adaptation.
5 Discussions and Intriguing Insights
Sparse vs. low-rank PEFT. As shown in Figure 2, OFT and LoRA achieve parameter-efficiency through sparsity and low rank, respectively. This suggests an intriguing analogy between OFT and LoRA, as sparsity and low rank represent arguably two of the most widely studied and exploited structural properties in matrices. To further enhance the scalability of OFT, more structured sparsity should be exploited, e.g., butterfly factorization Liu et al. (2024). Moreover, similar to AdaLoRA Zhang et al. (2023c), the sparsity level in OFT can be conditioned on the task and layer. Compared to low-rank PEFT, sparse PEFT approaches like OFT remain relatively underexplored, leaving many interesting open problems for future investigation.
Sequential vs. parallel adaptation. As shown in Figure 3, OFT and LoRA exemplify two distinct adaptation strategies: sequential adaptation and parallel adaptation, respectively. This contrast is particularly intriguing, as it explains why sequential adaptation benefits from orthogonality, while parallel adaptation naturally aligns with low rank. Sequential adaptation offers great expressiveness but is also more susceptible to error propagation and distortion of the pretrained model’s spectral properties. Enforcing orthogonality on $\bm{R}$ is therefore a natural choice, as it preserves these properties and helps prevent the accumulation of errors. Sparsity is the natural choice if we want to save parameters in orthogonal matrices. Parallel adaptation adds the adapter $\bm{R}$ to the pretrained model. In this case, we want $\bm{R}$ to be a dense update while maintaining parameter efficiency–a goal naturally achieved through low-rank matrices. This perspective may inspire new directions in adapter design.
Efficient orthogonality parameterization. OFT also highlights the importance of efficient parameterization of orthogonal matrices. In fact, the efficiency is closely tied to two factors: (1) the degree to which orthogonality needs to be approximated, and (2) the size of the set of orthogonal matrices considered. Our experiments indicate that exact orthogonality and the full orthogonal group are not strictly necessary, as parameterizations from the special orthogonal group and approximate orthogonality perform quite well in practice. This raises an open question: can we find even more efficient parameterizations with comparable performance?
Figure 4: Results of GPU memory usage for the same finetuning task. (a) OFT, LoRA and OFTv2 on Qwen2.5; (b) QLoRA and QOFT on NF4-quantized Qwen2.5; (c) QLoRA and QOFT on AWQ-quantized Qwen2.5.
6 Experiments on Scalability
Our experiments systematically evaluate OFTv2 along two key dimensions: (1) its scalability improvements over the original OFT, and (2) its finetuning performance across a diverse set of tasks from multiple domains. For both aspects, we compare OFTv2 and QOFT against the well-established, memory- and compute-efficient low-rank adaptation methods LoRA Hu et al. (2022b) and QLoRA Dettmers et al. (2023).
6.1 GPU Memory Efficiency
As depicted in Figure 1, OFTv2 achieves a $3\times$ reduction in GPU memory consumption compared to the original OFT when finetuning the Qwen2.5-7B model. Furthermore, QOFT significantly reduces memory consumption by enabling the orthogonal finetuning of quantized base models. In the following ablation studies comparing against both LoRA and QLoRA baselines, where QLoRA broadly refers to low-rank adaptation of quantized models without being limited to NormalFloat 4-bit quantization, we evaluate the actual GPU memory consumption during finetuning of Qwen2.5 models from 0.5B to 72B parameters. For a comprehensive analysis, we additionally incorporate the widely adopted quantization method AWQ Lin et al. (2024) for activation-aware quantization. The results are summarized in Figure 4. Our experimental results demonstrate that OFTv2 and QOFT achieve memory efficiency comparable to low-rank adaptation methods, with a consistent performance across model scales and data formats.
6.2 Computational Efficiency
[…]
Model Size GPUs LoRA OFTv2
[…]
Table 1: Training time (clock time) comparison: OFTv2 vs. LoRA on GSM8K for mathematical reasoning.
7 Experiments on Performance
Having established that OFTv2 achieves comparable memory and computational efficiency to low-rank adaptation methods, we then test its performance on a variety of tasks.
7.1 Encoder-Decoder Model: BART
We evaluate the finetuning of BART-large Lewis et al. (2019) on the XSum Narayan et al. (2018) and CNN/DailyMail Hermann et al. (2015) datasets for text summarization, reporting ROUGE-1/2/L scores for LoRA and OFTv2 under both full-precision and NormalFloat4 4-bit quantization. We further investigate different configurations by increasing the rank $r$ for LoRA and the block size $b$ for OFTv2. The results from these finetuning tasks are reported in Table 3. We observe that OFTv2/QOFT consistently outperforms LoRA/QLoRA across all tested configurations, while notably utilizing 47–53% fewer trainable parameters. The performance gain gets more obvious with increasing model capacity: at the maximum parameter budget, QOFT outperforms QLoRA by +0.93 ROUGE-1 on XSum (44.16 vs. 43.23), suggesting a more effective utilization of expanded adapters. Furthermore, the finetuning performance of OFTv2/QOFT further improves with an increase budget of trainable parameters.
Model Size GPUs QLoRA QOFT
[…]
Table 2: Clock time comparison of QOFT and QLoRA on OpenR1-Math-220k for mathematical reasoning.
Figure 5: Qualitative results from Dreambooth finetuning of Stable Diffusion 3.5 Large (8.1B parameters), with peak allocated GPU memory: LoRA (52.33 GB), OFT (52.32 GB), QLoRA (41.60 GB) and QOFT (41.53 GB).
Quant. LoRA / QLoRA OFTv2 / QOFT
# $uparrow # Params XSum\uparrow CNN/DailyMail\uparrow$
Full Prec. 4.33M 43.33 / 20.06 / 35.11 43.11 / 20.22 / 29.69 2.03M 43.36 / 20.21 / 35.31 43.27 / 20.29 / 29.71
8.65M 43.47 / 20.19 / 35.21 43.20 / 20.31 / 29.71 4.19M 43.85 / 20.69 / 35.83 43.72 / 20.73 / 30.22
17.30M 43.38 / 20.20 / 35.25 43.17 / 20.31 / 29.72 8.52M 44.12 / 20.96 / 36.01 44.08 / 21.02 / 30.68
NF4 4.33M 43.09 / 19.82 / 34.92 43.17 / 20.25 / 29.66 2.03M 43.10 / 19.92 / 35.00 43.31 / 20.37 / 29.74
8.65M 43.15 / 19.80 / 34.92 43.10 / 20.24 / 29.65 4.19M 43.72 / 20.58 / 35.68 43.71 / 20.74 / 30.22
17.30M 43.23 / 19.92 / 35.10 43.11 / 20.23 / 29.63 8.52M 44.16 / 20.98 / 36.09 44.10 / 21.05 / 30.69
Table 3: ROUGE-1, ROUGE-2, and ROUGE-L scores for BART-large finetuned on XSum and CNN/DailyMail.
Model Metric 16-bit 4-bit
LoRA OFTv2 QLoRA QOFT
7B # Params 39.98M 17.65M 39.98M 17.65M
[…]
13B # Params 62.59M 27.62M 62.59M 27.62M
[…]
Table 4: Finetuning results of Llama-2 models on WikiText-2 (perplexity) and GSM8K (test accuracy).
Model Type # Params AMC23 AQUA CMATH GaoKao Minerva Olympiad/ SAT
2023 En Math Bench Math
Qwen2.5-1.5B-it Baseline - 17.5 49.2 65.2 36.4 9.6 12.0 59.4
QLoRA 18.46M 15.0 42.5 61.5 29.6 8.1 8.9 59.4
QOFT 7.89M 27.5 53.1 68.5 41.0 11.8 14.4 81.2
Qwen2.5-1.5B Baseline - 0.0 18.9 4.0 4.2 2.6 2.4 28.1
QLoRA 18.46M 15.0 37.4 64.2 26.8 8.5 6.8 62.5
QOFT 7.89M 22.5 53.1 56.3 36.1 8.5 12.7 87.5
Qwen2.5-7B-it Baseline - 50.0 16.5 89.3 61.8 33.5 36.6 53.1
QLoRA 40.37M 30.0 48.0 88.8 50.1 25.4 19.7 68.8
QOFT 17.55M 52.5 70.9 90.5 63.6 33.5 37.6 96.9
Qwen2.5-7B Baseline - 25.0 55.1 61.2 42.9 11.8 29.9 71.9
QLoRA 40.37M 35.0 48.8 73.7 49.9 18.8 18.5 62.5
QOFT 17.55M 52.5 59.4 80.7 55.6 21.7 34.7 87.5
Qwen2.5-32B-it Baseline - 62.5 18.5 92.5 70.1 41.5 44.4 65.6
QLoRA 134.22M 62.5 71.7 94.0 71.2 39.7 46.8 96.9
QOFT 57.90M 75.0 83.1 94.7 73.5 41.5 48.7 100.0
Qwen2.5-32B Baseline - 35.0 23.2 35.7 46.8 20.2 25.2 62.5
QLoRA 134.22M 40.0 52.4 90.5 61.0 32.0 29.8 65.6
QOFT 57.90M 70.0 68.5 90.7 71.4 36.0 44.9 93.8
Table 5: Pass@1 performance of the Qwen2.5 series LLMs and its QLoRA/QOFT finetuned variants using the chain-of-thought reasoning distilled from DeepSeek R1.
7.2 Decoder-only Model: Llama-2 Series
We finetune Llama-2 7B and 13B models on the NLG datasets GSM8K Cobbe et al. (2021) and WikiText-2 Merity et al. (2017). To ensure fairness, we use the same set of hyperparameters for each method across datasets, precisions, and model scales. Both LoRA and QLoRA set rank to $r=16$. Both OFTv2 and QOFT set block size to $b=32$. Table 4 shows that OFTv2 consistently outperforms the low-rank adapter across different settings.
7.3 Decoder-only Model: Qwen2.5 Series
We perform supervised finetuning on the Huggingface OpenR1-Math-220k OpenR1-Team (2025) dataset—a large-scale mathematical reasoning corpus containing challenging problems and two to four reasoning traces distilled from DeepSeek R1 Guo et al. (2025). Following the evaluation protocol of Qwen2.5-Math Yang et al. (2024a), we report pass@1 performance on established math benchmarks: CMATH Wei et al. (2023), AMC23 Project-Numina , AQUA Ling et al. (2017), Olympiad Bench He et al. (2024), Gaokao 2023 En Liao et al. (2024), and Minerva Math Lewkowycz et al. (2022). Finetuning was only performed on NormalFloat 4-bit quantized base models due to the substantial memory requirements imposed by the large context window size (16384), necessary for training on a reasoning dataset. The results are reported in Table 5. The baseline method refers to the pre-trained Qwen2.5 models without any continual training. We observe that QOFT consistently outperforms both QLoRA and the base model across all evaluated scales and tasks, despite using significantly fewer trainable parameters. For instance, on the Qwen2.5-7B instruction-tuned model, QOFT achieves a 96.9% SAT Math accuracy compared to QLoRA’s 68.8%, while utilizing only 17.55M parameters (57% fewer than QLoRA’s 40.37M). This advantage scales robustly: the Qwen2.5-32B variant finetuned with QOFT attains 100% SAT Math accuracy, surpassing both the baseline (65.6%) and QLoRA (96.9%). These gains persist across mathematical reasoning tasks (e.g., 70.0% on AMC23 for QOFT-32B vs. QLoRA’s 40.0%), suggesting that orthogonal adaptation in quantized space better preserves the model’s reasoning capabilities compared to low-rank adaptation. The results demonstrate QOFT’s dual strength: parameter efficiency without sacrificing task performance, particularly in the quantized setting. In contrast, QLoRA-finetuned models can exhibit training instabilities Li et al. (2023), leading to model collapse where their performance fell below the base model. Appendix C gives more results on finetuning math-specific Qwen2.5 models.
7.4 Text-to-image Generative Models: SD-3.5
To assay the generality of the proposed methods across modalities, we perform Dreambooth Ruiz et al. (2023) finetuning on the latest Stable Diffusion 3.5 models Esser et al. (2024). Dreambooth finetunes text-to-image models using a limited set of images depicting the same subject. This process binds the subject to a unique token identifier, enabling subject-driven generation where the model synthesizes this subject in novel scenes beyond the training data. Qualitative results are shown in Figure 5 and Appendix D. We also report the actual peak GPU memory usage during the finetuning process in Appendix D. For finetuning the NormalFloat 4-bit quantized Stable Diffusion 3.5 Large model, QOFT requires slightly less GPU memory ($35.02$ GB) than the QLoRA method ($35.03$ GiB).
8 Concluding Remarks
OFTv2 advances orthogonal finetuning through three key innovations: (i) an input-centric reformulation using matrix–vector products, reducing training time by over 10× and peak memory by  3× without loss in performance; (ii) a Neumann series based approximation of the Cayley transform, improving numerical stability while preserving approximate orthogonality; and (iii) an extension to quantized models, which matches or surpasses QLoRA in speed, stability, and memory efficiency. Across BART, LLaMA2, Qwen2.5, and Stable Diffusion3.5 (0.5B–72B), OFTv2 achieves competitive performance with roughly half the trainable parameters and consistent memory savings.
9 Limitations
OFTv2 substantially improves upon OFT in both memory and computational efficiency, matching low-rank methods in memory usage across data types and training speed in the quantized setting. However, its full-precision fine-tuning remains slower. This limitation arises from fundamental differences: low-rank can be naturally maintained efficiently through two simple linear layers, while preserving orthogonality presents a greater optimization challenge. Additionally, low-rank approaches benefit from extensive community-driven engineering and optimization. Bridging this computational gap presents an interesting research direction.
Acknowledgment
The authors would like to sincerely thank Tim Z. Xiao, Le Chen, Yao Feng and Zhen Liu for suggestions and helpful discussions. The core idea was proposed by WL and ZQ, the experiments were conducted by ZQ, and the project was led and supervised by WL. The paper was drafted by WL and ZQ, and later polished by AW and BS.
Appendix
Appendix A Experimental Details
This section outlines the specifics of our experimental setup, including the optimizer, code frameworks, computational resources, evaluation methods, and detailed hyperparameters used for each experiment.
Training details.
We employed the Adam optimizer Kingma and Ba (2015) for all our training runs. The specific hyperparameters used for each experiment are detailed in the tables referenced below. These include learning rates, batch sizes, number of training epochs, and method-specific configurations: the rank $r$ for LoRA-based methods and the block size $b$ for OFTv2/QOFT. If not explicitly specified, the $r$ for LoRA-based methods is 16 and the block size $b$ for OFTv2/QOFT is set as 32. For the Wikitext dataset, hyperparameters are listed in Table 8. For the GSM8K dataset, hyperparameters are listed in Table 9. For the XSum dataset, hyperparameters are listed in Table 6. For the CNN/DailyMail dataset, hyperparameters are listed in Table 7. Since it is known that merging QLoRA adapter weights to its quantized base models leads to performance degradation [^1] Comparison of merging methods: https://kaitchup.substack.com/p/lora-adapters-when-a-naive-merge and distorts the real performance, for every experiment, we evaluate the fine-tuned model without merging the trainable parameters, but load them as extra adapter layers.
Hyperparameter LoRA OFTv2
BF16 NF4 BF16 NF4
$r=8$ $r=16$ $r=32$ $r=8$ $r=16$ $r=32$ $b=16$ $b=32$ $b=64$ $b=16$ $b=32$ $b=64$
Learning rate 1e-4 1e-4 1e-4 1e-4 1e-4 1e-4 4e-4 4e-4 4e-4 4e-4 4e-4 4e-4
Epoch 10 10 10 10 10 10 5 5 5 5 5 5
Batch size 32 32 32 32 32 32 32 32 32 32 32 32
Gradient Accumulation 4 4 4 4 4 4 4 4 4 4 4 4
Table 6: Hyper-parameter setup of fine-tuning BART-large on XSum with LoRA and OFTv2.
Hyperparameter LoRA OFTv2
BF16 NF4 BF16 NF4
$r=8$ $r=16$ $r=32$ $r=8$ $r=16$ $r=32$ $b=16$ $b=32$ $b=64$ $b=16$ $b=32$ $b=64$
Learning rate 1e-4 1e-4 1e-4 1e-4 1e-4 1e-4 4e-4 4e-4 4e-4 4e-4 4e-4 4e-4
Epoch 5 5 5 5 5 5 5 5 5 5 5 5
Batch size 64 64 64 64 64 64 64 64 64 64 64 64
Gradient Accumulation 4 4 4 4 4 4 4 4 4 4 4 4
Table 7: Hyper-parameter setup of fine-tuning BART-large on CNN/DailyMail with LoRA and OFTv2.
Hyperparameter LoRA OFTv2
BF16 NF4 BF16 NF4
7B 13B 7B 13B 7B 13B 7B 13B
Learning rate 2e-4 2e-4 2e-4 2e-4 2e-4 2e-4 2e-4 2e-4
Epoch 10 10 10 10 10 10 10 10
Batch size 16 16 16 16 16 16 16 16
Gradient Accumulation 2 2 2 2 2 2 2 2
Table 8: Hyper-parameter setup of fine-tuning Llama 2 on Wikitext-2 with LoRA and OFTv2.
Hyperparameter LoRA OFTv2
BF16 NF4 BF16 NF4
7B 13B 7B 13B 7B 13B 7B 13B
Learning rate 2e-4 2e-4 2e-4 2e-4 8e-4 8e-4 8e-4 8e-4
Epoch 10 10 10 10 10 10 10 10
Batch size 16 16 16 16 16 16 16 16
Gradient Accumulation 4 4 4 4 4 4 4 4
Table 9: Hyper-parameter setup of fine-tuning Llama 2 on GSM8K with LoRA and OFTv2.
Code framework.
Our method is implemented using the Hugging Face PEFT [^2] https://huggingface.co/docs/peft/en/index framework, a widely adopted open-source framework providing state-of-the-art parameter-efficient fine-tuning of pre-trained large language models and diffusion models. The implementation of OFTv2 will be released on Hugging Face PEFT soon, to allow for easy reproduction of our training results. We utilized the Hugging Face TRL library for supervised fine-tuning [^3] https://github.com/huggingface/trl. For the base model quantization, we leveraged bitsandbytes [^4] https://github.com/bitsandbytes-foundation/bitsandbytes for the NormalFloat 4-bit quantization and the QLoRA finetuning, and AutoAWQ [^5] https://github.com/casper-hansen/AutoAWQ for AWQ quantization.
Pretrained models.
Our work utilized several pre-trained large language models. Specifically, we employed models from the Qwen2.5 model series [^6] https://huggingface.co/collections/Qwen/qwen25-66e81a666513e518adb90d9e, which are available under the permissive Apache 2.0 license. We also leveraged the Llama 2 models [^7] https://huggingface.co/collections/meta-llama/metas-llama2-models-675bfd70e574a62dd0e40541, governed by the Llama 2 license. Additionally, for the text summarization tasks, the BART-large model was used, which is also distributed under the Apache 2.0 license. For the text-to-image generation, we utilized the Stable Diffusion 3.5 models, which are under the Stability AI Community license. We have adhered to all respective licensing agreements for these models throughout our work.
Dataset.
The experiments in this study utilized a diverse range of publicly available datasets to ensure comprehensive evaluation. For finetuning language modeling tasks, we employed the Wikitext-2888https://huggingface.co/datasets/Salesforce/wikitext dataset, which is distributed under the CC-BY-SA-3.0 license. Text summarization performance was assessed by fine-tuning on the CNN / DailyMail Dataset [^9] https://huggingface.co/datasets/abisee/cnn_dailymail, also licensed under Apache 2.0, and the XSum dataset [^10] https://huggingface.co/datasets/EdinburghNLP/xsum, which is available under the MIT license. For finetuning mathematical reasoning capabilities, we used the GSM8K [^11] https://huggingface.co/datasets/openai/gsm8k dataset, available under the MIT license, and the OpenR1-Math-220k [^12] https://huggingface.co/datasets/open-r1/OpenR1-Math-220k dataset, which can be used under the Apache 2.0 license. The Dreambooth dataset [^13] https://huggingface.co/datasets/google/dreambooth for fine-tuning the diffusion models are under the cc-by-4.0 license.
Compute Resources.
All the training tasks are performed on a NVIDIA HGX H100 8-GPU System node with 80GB memory each. We used a single NVIDIA H100 NVL GPU with 94GB memory to benchmark the memory usage.
Appendix B Effect of Neumann Series Terms in Orthogonal Parameterization
OFTv2 employs the Cayley-Neumann parameterization to improve the training efficiency; the number of Neumann series terms becomes a hyperparameter. We conducted an additional ablation study to evaluate the impact of the number of Neumann series terms on finetuning performance for WikiText. The results are reported in Table 10. We observe that when the number of Neumann terms is too small (e.g., 2), the approximation error to orthogonality slightly degrades performance. For the experiments reported in the main paper, we used five Neumann terms, which we found to be well-suited across all evaluated tasks.
Model Method 2 terms 3 terms 4 terms 5 terms 6 terms
Llama 2 7B OFTv2 6.22 6.15 6.14 6.13 6.14
Llama 2 13B OFTv2 5.11 5.00 4.99 4.98 4.99
Llama 2 7B QOFT 5.70 5.62 5.58 5.60 5.61
Llama 2 13B QOFT 5.14 5.02 5.04 5.05 5.05
Table 10: Effect of Neumann Series Terms on the Llama-2 Models
Appendix C Mathematical Reasoning with Qwen2.5
Training details.
We fine-tuned the Qwen2.5 models using QLoRA or QOFT on a random subset of 50,000 samples from the Huggingface OpenR1-Math-220k dataset OpenR1-Team (2025). For each method and benchmark, we selected the best-performing model after trying learning rates of $1\times 10^{-5}$, $2\times 10^{-5}$, $5\times 10^{-5}$, and $1\times 10^{-4}$. We used a batch size of 16 for the 1.5B models and 8 for the 7B and 32B models, with 2 gradient accumulation steps for all. A cosine learning rate scheduler was employed, with a minimum learning rate set to 10% of the initial value.
Evaluation details.
For evaluating the Qwen2.5 base models and the QLoRA or QOFT fine-tuned versions, we utilized the same evaluation pipeline as Qwen2.5-Math [^14] https://github.com/QwenLM/Qwen2.5-Math. This framework provides robust tools for parsing and evaluating mathematical expressions and problem-solving steps, ensuring accurate and consistent assessment of model performance on these mathematical benchmarks. More specifically, we report the model’s pass@1 performance, i.e., the performance on the first attempt for a given task, obtained by utilizing the Qwen2.5 Chain-of-Though question prompt (Figure 6).
[…]
Figure 6: Prompt template used for evaluating Qwen2.5 series models on mathematical reasoning benchmarks.
Model Method # Params AMC23 AQUA CMATH GaoKao Minerva Olympiad/ SAT
2023 En Math Bench Math
Qwen2.5-1.5B-math-it QLoRA 18.46M 27.5 33.5 86.8 43.6 15.4 15.1 46.9
QOFT 7.89M 45.0 70.9 87.2 60.5 25.4 32.0 93.8
Qwen2.5-1.5B-math QLoRA 18.46M 25.0 31.5 49.0 36.9 10.7 12.9 50.0
QOFT 7.89M 27.5 31.5 55.5 37.7 13.6 14.4 37.5
Qwen2.5-7B-math-it QLoRA 40.37M 32.5 34.6 89.8 47.0 18.8 18.2 53.1
QOFT 17.55M 52.5 76.8 92.7 66.8 35.7 41.6 93.8
Qwen2.5-7B-math QLoRA 40.37M 30.0 38.6 75.7 48.6 21.0 20.4 50.0
QOFT 17.55M 30.0 40.6 81.7 49.4 21.3 20.4 50.0
Table 11: The pass@1 performance of the Qwen2.5 series math-specific large language fine-tuned with QLoRA/QOFT by the Chain-of-Thought reasoning.
Appendix D Subject-driven Generation with Stable diffusion 3.5
Here we provide additional qualitative results of fine-tuning the Stable Diffusion 3.5 Medium model in Figure 7.
Figure 7: Qualitative results from Dreambooth fine-tuning of Stable Diffusion 3.5 Medium (8.1B parameters), with peak allocated GPU memory: LoRA (38.00 GB), OFT (38.02 GB), QLoRA (35.03 GB) and QOFT (35.02 GB).
The actual GPU memory usage during LoRA and OFTv2 fine-tuning is summarized in Table 12. As shown, OFTv2/QOFT demonstrates memory efficiency similar to LoRA and QLoRA, regardless of data precision or model scale.
SD 3.5 Medium SD 3.5 Large
LoRA 38.00 GB 52.33 GB
OFTv2 38.02 GB 52.32 GB
QLoRA 35.03 GB 41.60 GB
QOFT 35.02 GB 41.53 GB
Table 12: Actual GPU memory usage during fine-tuning: LoRA, QLoRA, OFTv2, and QOFT applied on Stable Diffusion 3.5 Medium and Large.